OpenStudy (anonymous):

find the integral of {1}/ (sqrt){1-x^2}

OpenStudy (anonymous):

well, just memorized that integral of 1/sqrt(1-x^2) is arcsin(x)

OpenStudy (anonymous):

To derive the formula for yourself, use a trig sub: x = sin(u) dx = cos(u) du So you have Int[ cos(u)/√( 1 - sin²(u) ) du ] = Int[ cos(u)/√( cos²(u) ) du ] = Int[ cos(u) / cos(u) du ] = Int[ du ] = ∙∙∙

OpenStudy (anonymous):

i will try that @SithsAndGiggles

OpenStudy (anonymous):

or just memorize it like every other one. i.e arcsec/arctan